On July 1, 2020, Shroff Company leased a warehouse building under a 10-year lease agreement. The lease requires quarterly lease payments of $6,000. The first lease payment is due on September 30, 2020. The lease was reported as a finance lease using an 8% annual interest rate.

a. Prepare the journal entry to record the commencement of the lease on July 1, 2020.
b. Prepare the journal entries that would be necessary on September 30 and December 31, 2020.
c. Post the entries from parts a and b in their appropriate T-accounts.
d. Prepare a financial statement effects template to show the effects template to show the effects of the entries from parts a and b on the balance sheet and income statement.

Note: Round answers to the nearest dollar. Use rounded answers for subsequent computations.

  • Journal entries
  • T-accounts
  • Financial statement effects template

Note:Use negative sign with your answers, when appropriate.

Balance Sheet Income Statement
Transactions Cash Asset Noncash
Assets
Contra
Assets
Liabilities Contributed
Capitals
Earned
Capital
Revenues Expenses Net
Income
7/1: Finance lease commences. Answer Answer Answer Answer Answer Answer Answer Answer Answer
9/30: Amortization on leased asset. Answer Answer Answer Answer Answer Answer Answer Answer Answer
9/30: Made quarterly lease payment. Answer Answer Answer Answer Answer Answer Answer Answer Answer
12/31: Amortization on leased asset. Answer Answer Answer Answer Answer Answer Answer Answer Answer
12/31: Made quarterly lease payment. Answer Answer Answer Answer Answer Answer Answer Answer Answer
Answer #1
a. Journal entries;
Date General journal Debit Credit
July 1,2020 Leased building 164133
lease payable 164133
(lease recorded)
Present value of lease payments;
Lease paid quarterly
Discount factor=Interest per quarter=8%*(3/12)=2%
Total quarterly lease payments=10*4=40 periods
Present value of lease payments=6000*Present value at 2% for 40 years=6000*27.35548=$ 164133
b. Date General journal Debit Credit
Sep 30. Depreciation expense (164133/10)*(3/12) 4103
Accumulated depreciation 4103
(Depreciation recorded for 3 months)
Sep 30. Interest expense (164133*8%*3/12) 3283
Lease payable (6000-3283) 2717
Cash 6000
(Lease payment made)
Dec 31. Depreciation expense (164133/10)*(3/12) 4103
Accumulated depreciation 4103
(Depreciation recorded for 3 months)
Dec 31. Interest expense (164133-2717)*8%*3/12 3228
Lease payable (6000-3228) 2772
Cash 6000
(Lease payment made)
c. Leased building Lease payable
Date Debit Date Credit Date Debit Date Credit
Jul 1. 164133 Sep 30. 2717 Jul 1. 164133
Dec 31. 2772
5489 164133
Bal. 158644
Accumulated depreciation Interest expense
Date Debit Date Credit Date Debit Date Credit
Sep 30. 4103 Sep 30. 3283
Dec 31. 4103 Dec 31. 3228
0 8206 6511 0
Bal. 8206 Bal. 6511
Depreciation expense
Date Debit Date Credit
Sep 30. 4103
Dec 31. 4103
8206 0
Bal. 8206
d. Financial statement effects template
Balance sheet Income statement
Transaction Cash asset Noncash asset Contra assets Liabilities Contributed capitals Earned capital Revenues Expenses Net income
7/1:Finance lease commences 164133 164133
9/30: Amortization on leased asset -4103 -4103 -4103 -4103
9/30: Made quarterly lease payment -6000 -2717 -3283 -3283 -3283
12/31: Amortization on leased asset -4103 -4103 -4103 -4103
12/31: Made quarterly lease payment -6000 -2772 -3228 -3228 -3228